OpenStudy (anonymous):

solve: square root of x =20

OpenStudy (anonymous):

-4.47 or 4.47

OpenStudy (anonymous):

19 are remaining

OpenStudy (anonymous):

thats not one of the answers :(

OpenStudy (anonymous):

give me the options

OpenStudy (anonymous):

\[\sqrt{x}=20\]

OpenStudy (anonymous):

thats the equation

OpenStudy (anonymous):

x=400

OpenStudy (anonymous):

next

OpenStudy (anonymous):

so it 400

OpenStudy (anonymous):

or \[\pm400\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

20 * 20

OpenStudy (anonymous):

next one is \[\sqrt{x}=16\]

OpenStudy (texaschic101):

16 * 16

OpenStudy (anonymous):

16*16=256

OpenStudy (anonymous):

next

OpenStudy (anonymous):

Which set of numbers could represent the lengths of the sides of a right triangle? 8, 10, 18 8, 10, 14 6, 8, 14 6, 8, 10 .

OpenStudy (texaschic101):

if you have the square root of x = a number, then the answer is that number ^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

6,8,10

OpenStudy (anonymous):

@xo_ellaaavasquez pythagorean theorem 36+64=100 then sqrt100=10 , then it's 6,8,10

OpenStudy (anonymous):

Which set of numbers could represent the lengths of the sides of a right triangle? \[7,\sqrt{47}, 9 \] \[ 5,\sqrt{39}, 8 \] \[ 3,\sqrt{29}, 6 \] \[9,\sqrt{10}, 11\]

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

solve \[\sqrt{3x-4}=3\]

OpenStudy (anonymous):

x=13/3 ,,,,next one

OpenStudy (anonymous):

solve \[5\sqrt{x}+3=7\]

OpenStudy (anonymous):

16/25

OpenStudy (anonymous):

next one

OpenStudy (anonymous):

Find the missing side length of the right triangle 

OpenStudy (anonymous):

9.79

OpenStudy (anonymous):

A.6 B.16 C.\[\sqrt{146}\] D.\[4\sqrt{6}\]

OpenStudy (anonymous):

those are the answers for that question find the missing length

OpenStudy (anonymous):

D. 46√

OpenStudy (anonymous):

c'mon next one

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Let y = 45,000(0.97)x represent the purchasing power of $45,000 with an inflation rate of 3% per year, where x represents the number of years. What is the purchasing power when x = 0?

OpenStudy (anonymous):

45000 i think because 0.97^0=1 ,then the answer i think is 45000

OpenStudy (anonymous):

Erica is in training for a cross country race. She has been training by running 3 miles each day. She will be increasing her running by this weekly formula:, where x represents the number of weeks and y represents the distance run. What is the value if x = 0? 0 1.1 3 3.3

OpenStudy (anonymous):

@xo_ellaaavasquez 3 i guess , next one

OpenStudy (anonymous):

Let y = 32,000(0.94)x represent the purchasing power of $32,000 with an inflation rate of 6% per year, where x represents the number of years. The table represents the values for the four years after the initial one. Which year has an incorrect value? Years later (x) 1 2 3 4 Purchasing Power $ (y) 30,080 28,275 26579 24,592

OpenStudy (anonymous):

D

OpenStudy (anonymous):

next

OpenStudy (anonymous):

The equation, y = 16,000(0.82)x, represents the value of a $16,000 vehicle after x years of age. Calculate the value for x = 8. Given that this represents the value of the vehicle 8 years later, does this value make sense? yes or no

OpenStudy (anonymous):

try this value 16000(0.82^8) = $3270.62

OpenStudy (anonymous):

so it does make sense because thats what the question is asking lol

OpenStudy (anonymous):

so it's correct ??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the old is gold hhhh

OpenStudy (anonymous):

The number of farms in a certain area can be represented by the equation, y = 85,000(0.95)x, where x represents the number of years after 1990. Predict the number of farms in the area in 2008 (x = 18), to the nearest whole number. 35,540 33,763 32,075 30,471

OpenStudy (anonymous):

y = 85000(0.95)^x then 85000(0.95^18)=33763

OpenStudy (anonymous):

Let y = 45,000(0.97)x represent the purchasing power of $45,000 with an inflation rate of 3% per year, where x represents the number of years. Predict the purchasing power in 10 years (x = 10), rounded to the nearest dollar. $45,000 $41,256 $38,528 $33,184

OpenStudy (anonymous):

would it be B

OpenStudy (anonymous):

D

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

Suppose that y = 37,000(1.005)x represents the median household income in the U.S. , where x represent the number of years after 1980. Predict the median household income in 2010 (x = 30), rounded to the nearest dollar. $111,215 $82,523 $42,972 $52,900

OpenStudy (anonymous):

and yes you are

OpenStudy (anonymous):

$42,972

OpenStudy (anonymous):

i have a couple questions left omg how can i freaking thank you aghhh thanks super much :,)

OpenStudy (anonymous):

no problemo

OpenStudy (anonymous):

This is the table for the billions of dollars in book sales for the first four years after 1990, using the equation . They have been rounded to the nearest whole number. Which of the years has an incorrect value? Years after 1990 (x) 1 2 3 4 Billions of dollars (y) 20 21 22 23 A. 1 B. 2 C. 3 D. 4 @kyusakazaki

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

wait a little bit

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think it's number 4 check it

OpenStudy (anonymous):

its correct

OpenStudy (anonymous):

23.7 near to 24 then it's correct

OpenStudy (anonymous):

yes well thank you so much helping ima get some sleep thank you so much

OpenStudy (anonymous):

remain one question

OpenStudy (anonymous):

??

OpenStudy (anonymous):

sorry i did a miscalculation